OpenStudy (anonymous):

Find the inverse of the function. f(x) = x3 - 3

OpenStudy (campbell_st):

start with \[y = x^3 - 3\] swap x and y \[x = y^3 - 3\] make y the subject add 3 to both sides \[x + 3 = y^3\] can you finish the last step..?

OpenStudy (anonymous):



OpenStudy (campbell_st):

yes \[y = \sqrt[3]{x + 3}\]